Ticket Number: HEA-NOC/20020415-04 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: Global Crossing
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020415 14:50 UTC +1
Ticket Closed: Problem End:
Lines to Abilene, CAnet, Star Tap, teleglobe and Geant are down.
All traffic should be rerouting through UUnet, but with impeeded performance.
20020415 15:00 BN Contacting Global Crossing.
20020415 15:25 mn GC advise of cut in wet cable near Bude this
morning, compounded by failure near Kilmore Quay at 14:50. Engineer
despatched to Kilmore Quay to bring up one limb of SDH ring to
20020415 15:54 BN Faults reported to GC. Several tickets have been issued.
DUB-NY-VC4SU101 - 436270
DUB-NY-VC4SU102 - 436281
DUB-NY-VC4SU103 - 436287
DUBLIN/HEA-LONDON/HEA-VC4SU101 - 436294
DUBLIN/HEA-FRANKFURT/HEA-VC4SU102 - 436300
20020415 15:55 mn GC report fault located ~24km inland from Kilmore Quay.
20020415 17:44 BN GC attempting to migrate us onto operation fibre, HEAnet confirm lines are still down.
20020415 18:15 BN Circuits come back up. Traffic starts to flow. GC confirm we have been migrated to working fibre. SDH ring not yet in place so further outages possible. Also, as work will be taking place on these circuits some very brief outages are expected this evening before normal service is fully resumed. GC will contact us morning of 20020416 to check up.
20020416 09:11 BN Brief outages recorded between 19:20 and 19:24 on 20020415. These outages are in line with the possible outages outlined in the previous ticket update.
Time to Fix:
3hrs 30mins for initial restoration of circuits.